Numerical aperture of a fibre is 0.5 and core refractive index is 1.48. Find cladding refractive index and acceptance angle .

1
Expert's answer
2022-03-06T15:16:31-0500

"\beta=\sqrt{n_2^2-n_1^2}=1.09,"

"\alpha=\arctan 0.5=26.6°."